OpenStudy (anonymous):

If m/n=5/6, then what is 3m+2n? (a)0 (b)2 (c)7 (d)10 (e)Cannot be determined from the information given

OpenStudy (anonymous):

let m=5x, and n=6x. then 3m+2n = 3(5x)+2(6x)=15x+12x= 27x. Depending on what x is it can be any of the answers above--> correct answer in this context is e.